OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. R(4, 0), m = 5

OpenStudy (amistre64):

standard form is: Ax + By = C such that A is not negative

OpenStudy (amistre64):

this can be reconstructed from the point slope form: given a point (p,q) and a slope (m) y = m(x-p) +q

OpenStudy (amistre64):

oh, and A,B are integers as well

OpenStudy (anonymous):

@amistre64 so far I have 5x-4+0=y . I can't seem to figure it out

OpenStudy (amistre64):

y = 5(x-4)+0 y = 5x - 20 20 + y= 5x 20 = 5x - y 5x - y = 20

OpenStudy (anonymous):

ah! thank you! I must've misprinted when I wrote it out

OpenStudy (amistre64):

it happens :)